# Trust in Autonomous Human–Robot Collaboration: Effects of Responsive Interaction Policies

### v1

@@ Abstract
Trust plays a central role in human–robot collaboration, yet its formation is rarely examined under the constraints of fully autonomous interaction. This pilot study investigated how interaction policy influences trust during in-person collaboration with a social robot operating without Wizard-of-Oz control or scripted repair. Participants completed a multi-stage collaborative task with a mobile robot that autonomously managed spoken-language dialogue, affect inference, and task progression. Two interaction policies were compared: a responsive policy, in which the robot proactively adapted its dialogue and assistance based on inferred interaction state, and a neutral, reactive policy, in which the robot provided only direct, task-relevant responses when prompted. Responsive interaction was associated with significantly higher post-interaction trust under viable communication conditions, despite no reliable differences in overall task accuracy. Sensitivity analyses indicated that affective and experiential components of trust were more sensitive to communication breakdown than evaluative judgments of reliability, and that as language-mediated interaction degraded, the trust advantage associated with responsiveness attenuated and ratings became less clearly interpretable as calibrated evaluations of collaborative competence. These findings suggest that trust in autonomous human–robot interaction emerges from process-level interaction dynamics and operates within constraints imposed by communication viability, highlighting the importance of evaluating trust under real autonomy conditions when designing interactive robotic systems.
Keywords human-robot collaboration • HRI • socially assistive robotics • cobots • autonomous robot systems • embodied AI • spoken language interaction • trust in human-robot interaction • affect-adaptive robot systems
@@ 1 Introduction
As artificial intelligence (AI) technologies advance, they are increasingly integrated into robotic systems, enabling more adaptive, autonomous, and context-sensitive behaviour in real-world environments. This convergence has accelerated the deployment of robots across safety-critical domains such as manufacturing [11], mining [18], and healthcare [12], where robots are now expected to operate alongside humans rather than in isolation [15, 37]. In these collaborative settings, successful deployment depends not only on technical performance and safety guarantees, but also on whether human users are willing to rely on, communicate with, and coordinate their actions around robotic partners [8, 16]. Trust has therefore emerged as a central determinant of adoption and effective use in human–robot collaboration (HRC) [41, 8]. Insufficient trust can lead to disuse or rejection of automation, while excessive trust risks overreliance and accidents—particularly in environments characterized by uncertainty or incomplete information [13].
@@
A substantial body of human–robot interaction (HRI) research has examined how robot behaviour influences user trust, perceived reliability, and cooperation across industrial and social contexts [35, 17]. Trust is typically conceptualized as a multidimensional construct encompassing cognitive evaluations of competence, predictability, and reliability, alongside a behavioural willingness to collaborate toward shared goals under conditions of risk or uncertainty [30, 21, 13]. Despite this multidimensional framing, empirical studies have most often operationalized trust using post-interaction self-report questionnaires collected following short, highly controlled, and often scripted interactions. While such measures provide valuable global assessments of user attitudes, they offer limited insight into how trust is negotiated, disrupted, and repaired during ongoing interaction—particularly in autonomous systems where errors and ambiguities are unavoidable [26].
@@
Recent advances in AI have expanded the range of interaction strategies available to autonomous robotic platforms in practice, enabling systems to move beyond fixed, scripted behaviours toward adaptive interaction policies that respond to user state and task context in real time [2]. Improvements in spoken-language processing, dialogue state tracking, and large language model–based reasoning now allow robots to adjust not only what they say, but when and how assistance is provided during collaboration [39]. In parallel, advances in affect inference from language and interaction cues have made it increasingly feasible for robots to incorporate estimates of user emotional state into interaction management [27, 37]. As a result, responsiveness in contemporary HRI is increasingly understood as a property of an underlying interaction policy, governing how a robot interprets cues, initiates support, and manages uncertainty, rather than as a surface-level social behaviour [5, 35, 17].
From an engineering perspective, responsiveness in autonomous human–robot interaction is not implemented as a single behavioural rule or surface-level cue, but emerges from the coordination of multiple system components responsible for perception, interpretation, and action [1]. In practice, responsive behaviour depends on the integration of dialogue management, state tracking, and inference mechanisms that estimate task progress, interaction uncertainty, and user affect. Proactive assistance based on this integrated state representation differs fundamentally from reactive, request-based behaviour; for example, a robot may offer clarification, encouragement, or pacing adjustments when confusion, hesitation, or frustration is inferred, rather than waiting for an explicit request for help [5].
@@
The present pilot study addresses these gaps by examining trust and collaboration during in-person interaction with a fully autonomous social robot in which participants collaborated with one of two versions of the same robot platform during a dialogue-driven puzzle task requiring shared problem solving. In both conditions, all interaction management—including speech recognition, dialogue state tracking, task progression, and response generation was handled and logged autonomously by a centralized dialogue agent without human intervention. In the responsive condition, the robot employed a proactive, affect-aware interaction policy, adapting its assistance based on conversational cues and inferred user affect (e.g., frustration or engagement). In the neutral condition, the robot followed a reactive policy, providing basic guidance and periodic check-ins, otherwise providing assistance only when explicitly requested.
This pilot study had three primary objectives: (1) to design and evaluate the feasibility of an autonomous spoken-language interaction system with affect-responsive behaviour on a mobile robot platform; (2) to assess whether a responsive interaction policy influences post-interaction trust and collaborative experience under autonomous conditions; and (3) to explore how behavioural and interaction-level indicators align with subjective trust evaluations.
@@ 2 Methods
This study employed a between-subjects experimental design to examine how robot interaction policy influences trust and collaboration during fully autonomous, in-person human–robot interaction. The sole experimental factor was the robot’s interaction policy. Participants were initially randomly assigned to condition; however, final condition allocation was partially constrained by participant attrition and scheduling, as described below.
@@ 2.2 Collaborative Task Design
The task structure was modelled after [25] and designed to examine collaboration under two distinct dependency conditions: enforced collaboration, in which successful task completion required the robot’s involvement, and optional collaboration, in which participants could choose whether and how to engage the robot. To operationalize these conditions in a naturalistic way, we embedded both tasks within a single immersive, narrative-driven puzzle scenario.
@@
Interactions took place in a shared physical workspace that included the Misty-II social robot and a participant-facing computer interface [19]. The interface displayed task materials, recorded participant responses, and supported progression through the game (see Figure 2). Importantly, it did not function as a control mechanism for the robot. Instead, the robot autonomously monitored task progression and participant inputs via the interface, adapting its dialogue and behaviour accordingly.
@@ 2.2.1 Task 1: Robot-Dependent Collaborative Reasoning
In the first task, participants were asked to identify who had taken the missing robot from a 6 × 4 grid of 24 potential “suspects.” Within the narrative, Misty informed participants that she had access to internal security footage but was unable to directly reveal the suspect’s identity due to laboratory security protocols. Instead, she could answer only yes/no questions about observable features of the individual in the footage (e.g., “Was the suspect wearing a hat?”). This constraint established the robot as the sole source of task-relevant information while maintaining narrative plausibility.
The suspect grid was displayed on the interface and allowed participants to click on and grey out eliminated candidates, while all questions were posed verbally to the robot (see Figure 2).
Figure 2: Task 1 interface including the 6 × 4 grid of 24 candidates. Participants could track those eliminated by clicking on subjects which would grey them out. A box was provided to input their final answer and a button included to move to the next task.
@@ 2.2.2 Task 2: Open-Ended Collaborative Problem Solving
The second task involved a more open-ended reasoning challenge designed to shift the dependency structure from enforced to optional collaboration. Participants were presented with multiple technical logs through a simulated terminal interface that could be used to infer the location of the missing robot (see Figure 3). The task was intentionally cryptic and difficult to solve within the allotted ten minutes without synthesizing information across multiple sources.
Figure 3: The task 2 interface presented multiple technical logs through a simulated terminal interface that could be used to determine the location of the missing robot.
@@
2.3 Study Protocol
Participants signed up for the study and completed a pre-session questionnaire before their in-person session via Qualtrics. The elapsed time between sign-up and the in-person session was at most week and at minimum immediately before the session. The pre-session questionnaire colleced basic demographics information and assessed baseline characteristics, including the Negative Attitudes Toward Robots Scale (NARS) and the short form of the Need for Cognition scale (NFC-s) [32, 7]. These measures were used to capture a baseline of individual differences that may moderate responses to robot interaction.
Upon arrival, participants were greeted by the researcher, provided with a brief overview of the session, and given instructions for effective communication with the robot. Once participants indicated readiness, the researcher exited the room, leaving the participant and robot to complete the interaction without human presence or observation. Participants initiated the interaction by clicking a start button on the interface and were informed that they could terminate the session at any time without penalty.
Following task completion, participants completed a 21-item post-interaction questionnaire assessing trust on a laptop. Participants then engaged in a brief debrief with the researcher and were awarded a \$15 gift card. Total session duration averaged approximately 30 minutes.
@@ 2.4.1 Self-Report Measures
Trust was assessed using two established self-report instruments commonly used in human–robot interaction research: the Trust Perception Scale–HRI (TPS-HRI) and the Trust in Industrial Human–Robot Collaboration scale (TI-HRC) [34, 9]. Both measures were adapted to reflect the specific dialogue-driven task context and interaction modality of the present study, while preserving the original constructs and response intent of each scale. 9 items were retained from the TI-HRC and 12 items from the TPS-HRI. Item wording was modified to reference the robot’s behaviour during the dialogue-driven collaborative tasks and response formats were adjusted to ensure interpretability for participants without prior robotics experience (see Appendix B for a full item list). Internal consistency of the adapted trust measures was evaluated within the present sample to assess measurement reliability under the modified wording and response formats. Given the pilot sample size, reliability estimates should be interpreted cautiously, but are reported to provide transparency regarding scale performance under fully autonomous interaction conditions.
All self-report scales demonstrated acceptable to strong internal consistency in the present sample. The Negative Attitudes Toward Robots scale (NARS) showed good reliability ($\alpha$ = .81), with subscale alphas ranging from .72 to .76. The Need for Cognition scale (NFC) demonstrated strong internal consistency ($\alpha$ = .87). Both post-interaction trust measures also exhibited high reliability. The Trust Perception Scale–HRI (TPS-HRI) showed strong internal consistency overall ($\alpha$ = .89), with subscale reliabilities ranging from .74 to .88. The Trust in Industrial Human–Robot Collaboration scale (TI-HRC) demonstrated very high internal consistency ($\alpha$ = .95).
@@ 2.4.2 Objective Measures
Table 3 outlines objective task metrics including the average number of dialogue turns, session duration, response time (robot dialogue latency), as well as the number of engaged versus frustrated responses the robot detected–all obtained from log data. Behavioural engagement metrics were derived from interaction logs and manually coded dialogue transcripts, including the proportion of turns characterized by communication breakdowns, human and robot reasoning, robot helpful and unhelpful guidance, robot and human affective engagement and other task-relevant human and robot contributions. See Appendix C for a full coding scheme.
@@ 2.5 Participants
A total of 29 participants were recruited from the Laurentian University community via word of mouth and the SONA recruitment system. Eligibility criteria required being 18 years or older, fluent in spoken and written English, and having normal or corrected-to-normal hearing and vision. Participants received a \$15 gift card as compensation for their time; some students additionally received partial credit for participating depending on their program of study. All procedures were approved by the Laurentian University Research Ethics Board (REB #6021966).
Although English fluency was an eligibility requirement, in-person observation during data collection revealed meaningful variability in participants’ functional spoken-language proficiency. The researcher therefore documented observed language proficiency and monitored interaction quality during each session in anticipation of potential speech-based system limitations.
Subsequent review of interaction transcripts and system logs indicated that a subset of sessions exhibited severe and sustained communication failure, characterized by fragmented or unintelligible automatic speech recognition (ASR) output and stalled dialogue. In these sessions, the robot was unable to extract sufficient linguistic content to maintain conversation, respond meaningfully to participant input, or support task progression. These interaction failures are described in detail in the Analytic Strategy and Results sections.
@@ 2.5.1 Randomization Check
Participants were initially assigned to the responsive or control condition by Qualtrics using an automated randomization procedure at the time of online sign-up. However, because several scheduled participants did not attend their in-person session, replacement participants were assigned to the next available session slot rather than re-randomized. As a result, the final condition assignment should be considered semi-random rather than strictly randomized.

### v2

# Trust in Autonomous Human-Robot Collaboration: Effects of Responsive Interaction Policies

@@ Abstract
Trust plays a central role in human-robot collaboration, yet its formation is rarely examined under the constraints of fully autonomous interaction. This pilot study investigated how interaction policy influences trust during in-person collaboration with a social robot operating without Wizard-of-Oz control or scripted repair. Participants completed a multi-stage collaborative task with a mobile robot that autonomously managed spoken-language dialogue, affect inference, and task progression. Two interaction policies were compared: a responsive policy, in which the robot proactively adapted its dialogue and assistance based on inferred interaction state, and a neutral, reactive policy, in which the robot provided only direct, task-relevant responses when prompted. Responsive interaction was associated with significantly higher post-interaction trust under viable communication conditions, despite no reliable differences in overall task accuracy. Sensitivity analyses indicated that affective and experiential components of trust were more sensitive to communication breakdown than evaluative judgments of reliability, and that as language-mediated interaction degraded, the trust advantage associated with responsiveness attenuated and ratings became less clearly interpretable as calibrated evaluations of collaborative competence. These findings suggest that trust in autonomous human-robot interaction emerges from process-level interaction dynamics and operates within constraints imposed by communication viability, highlighting the importance of evaluating trust under real autonomy conditions when designing interactive robotic systems.
Keywords Human-robot collaboration • Human-robot interaction • socially assistive robotics • autonomous robot systems • embodied AI • trust in human-robot interaction • affect-adaptive robot systems
@@ 1 Introduction
As artificial intelligence (AI) technologies advance, they are increasingly integrated into robotic systems, enabling more adaptive, autonomous, and context-sensitive behaviour in real-world environments. This convergence has accelerated the deployment of robots across safety-critical domains such as manufacturing [11], mining [18], and healthcare [12], where robots are now expected to operate alongside humans rather than in isolation [15, 37]. In these collaborative settings, successful deployment depends not only on technical performance and safety guarantees, but also on whether human users are willing to rely on, communicate with, and coordinate their actions around robotic partners [8, 16]. Trust has therefore emerged as a central determinant of adoption and effective use in human-robot collaboration (HRC) [41, 8]. Insufficient trust can lead to disuse or rejection of automation, while excessive trust risks overreliance and accidents—particularly in environments characterized by uncertainty or incomplete information [13].
@@
A substantial body of human-robot interaction (HRI) research has examined how robot behaviour influences user trust, perceived reliability, and cooperation across industrial and social contexts [35, 17]. Trust is typically conceptualized as a multidimensional construct encompassing cognitive evaluations of competence, predictability, and reliability, alongside a behavioural willingness to collaborate toward shared goals under conditions of risk or uncertainty [30, 21, 13]. Despite this multidimensional framing, empirical studies have most often operationalized trust using post-interaction self-report questionnaires collected following short, highly controlled, and often scripted interactions. While such measures provide valuable global assessments of user attitudes, they offer limited insight into how trust is negotiated, disrupted, and repaired during ongoing interaction—particularly in autonomous systems where errors and ambiguities are unavoidable [26].
@@
Recent advances in AI have expanded the range of interaction strategies available to autonomous robotic platforms in practice, enabling systems to move beyond fixed, scripted behaviours toward adaptive interaction policies that respond to user state and task context in real time [2]. Improvements in spoken-language processing, dialogue state tracking, and large language model-based reasoning now allow robots to adjust not only what they say, but when and how assistance is provided during collaboration [39]. In parallel, advances in affect inference from language and interaction cues have made it increasingly feasible for robots to incorporate estimates of user emotional state into interaction management [27, 37]. As a result, responsiveness in contemporary HRI is increasingly understood as a property of an underlying interaction policy, governing how a robot interprets cues, initiates support, and manages uncertainty, rather than as a surface-level social behaviour [5, 35, 17].
From an engineering perspective, responsiveness in autonomous human-robot interaction is not implemented as a single behavioural rule or surface-level cue, but emerges from the coordination of multiple system components responsible for perception, interpretation, and action [1]. In practice, responsive behaviour depends on the integration of dialogue management, state tracking, and inference mechanisms that estimate task progress, interaction uncertainty, and user affect. Proactive assistance based on this integrated state representation differs fundamentally from reactive, request-based behaviour; for example, a robot may offer clarification, encouragement, or pacing adjustments when confusion, hesitation, or frustration is inferred, rather than waiting for an explicit request for help [5].
@@
The present pilot study addresses these gaps by examining trust and collaboration during in-person interaction with a fully autonomous social robot in which participants collaborated with one of two versions of the same robot platform during a dialogue-driven puzzle task requiring shared problem solving. In both conditions, all interaction management, including speech recognition, dialogue state tracking, task progression, and response generation was handled and logged autonomously by a centralized dialogue agent without human intervention. In the responsive condition, the robot employed a proactive, affect-aware interaction policy, adapting its assistance based on conversational cues and inferred user affect (e.g., frustration or engagement). In the neutral condition, the robot followed a reactive policy, providing basic guidance and periodic check-ins, otherwise providing assistance only when explicitly requested.
The pilot study had three primary objectives: (1) to design and evaluate the feasibility of an autonomous spoken-language interaction system with affect-responsive behaviour on a mobile robot platform; (2) to assess whether a responsive interaction policy influences post-interaction trust and collaborative experience under autonomous conditions; and (3) to explore how behavioural and interaction-level indicators align with subjective trust evaluations.
@@ 2 Methods
This study employed a between-subjects experimental design to examine how robot interaction policy influences trust and collaboration during fully autonomous, in-person human-robot interaction. The sole experimental factor was the robot’s interaction policy. Participants were initially randomly assigned to condition; however, final condition allocation was partially constrained by participant attrition and scheduling, as described below.
@@
2.1 Study Protocol
Participants signed up for the study, provided informed consent, were randomly assigned to a group, and completed a pre-session questionnaire before their in-person session via a survey hosted on Qualtrics. The elapsed time between sign-up and the in-person session was at most a week and at minimum immediately before the session. The pre-session questionnaire collected basic demographics information and assessed baseline characteristics, including the Negative Attitudes Toward Robots Scale (NARS) and the short form of the Need for Cognition scale (NFC-s) [32, 7]. These measures were used to capture a baseline of individual differences that may moderate responses to robot interaction.
Upon arrival, participants were greeted by the researcher, provided with a brief overview of the session, and given instructions for effective communication with the robot. Once participants indicated readiness, the researcher exited the room, leaving the participant and robot to complete the interaction without human presence or observation (see Figure 7). Participants initiated the interaction by clicking a start button on the interface and were informed that they could terminate the session at any time without penalty.
Following task completion, participants completed a 21-item post-interaction questionnaire assessing trust on a provied laptop. Participants then engaged in a brief debrief with the researcher and were awarded a \$15 gift card. Total session duration averaged approximately 30 minutes.
@@ 2.3 Collaborative Task Design
The task structure was modelled after the interactive and immersive puzzle game design by [25] to examine collaboration under two distinct dependency conditions: enforced collaboration, in which successful task completion required the robot’s involvement, and optional collaboration, in which participants could choose whether and how to engage the robot. To operationalize these conditions in a naturalistic way, we embedded both tasks within a single immersive, narrative-driven puzzle scenario.
@@
Interactions took place in a shared physical workspace that included the Misty-II social robot and a participant-facing computer interface [19] at Laurentian University. The interface displayed task materials, recorded participant responses, and supported progression through the game. Importantly, it did not function as a control mechanism for the robot. Instead, the robot autonomously monitored task progression and participant inputs via the interface, adapting its dialogue and behaviour accordingly.
@@ 2.3.1 Task 1: Robot-Dependent Collaborative Reasoning
In the first task, participants were asked to identify who had taken the missing robot from a 6 × 4 grid of 24 potential “suspects.” Within the narrative, Misty informed participants that she had access to internal security footage but was unable to directly reveal the suspect’s identity due to laboratory security protocols. Instead, she could answer only yes/no questions about observable features of the individual in the footage (e.g., “Was the suspect wearing a hat?”). This constraint established the robot as the sole source of task-relevant information while maintaining narrative plausibility. The suspect grid was displayed on the interface and allowed participants to click on and grey out eliminated candidates, while all questions were posed verbally to the robot (see Figure 7).
@@
Figure 2: Task 1 interface including the 6 ×× 4 grid of 24 candidates. Participants could track those eliminated by clicking on subjects which would grey them out. A box was provided to input their final answer and a button included to move to the next task.
@@ 2.3.2 Task 2: Open-Ended Collaborative Problem Solving
The second task involved a more open-ended reasoning challenge designed to shift the dependency structure from enforced to optional collaboration. Participants were presented with multiple technical logs through a simulated terminal interface that could be used to infer the location of the missing robot. The task was intentionally cryptic and difficult to solve within the allotted ten minutes without synthesizing information across multiple sources.
@@ 2.4.1 Self-Report Measures
Trust was assessed using two established self-report instruments commonly used in human-robot interaction research: the Trust Perception Scale-HRI (TPS-HRI) and the Trust in Industrial Human-Robot Collaboration scale (TI-HRC) [34, 9]. Both measures were adapted to reflect the specific dialogue-driven task context and interaction modality of the present study, while preserving the original constructs and response intent of each scale. 9 items were retained from the TI-HRC and 12 items from the TPS-HRI. Item wording was modified to reference the robot’s behaviour during the dialogue-driven collaborative tasks and response formats were adjusted to ensure interpretability for participants without prior robotics experience (see Appendix B for a full item list). Internal consistency of the adapted trust measures was evaluated within the present sample to assess measurement reliability under the modified wording and response formats. Given the pilot sample size, reliability estimates should be interpreted cautiously, but are reported to provide transparency regarding scale performance under fully autonomous interaction conditions.
All self-report scales demonstrated acceptable to strong internal consistency in the present sample. The Negative Attitudes Toward Robots scale (NARS) showed good reliability ($\alpha$ = .81), with subscale alphas ranging from .72 to .76. The Need for Cognition scale (NFC) demonstrated strong internal consistency ($\alpha$ = .87). Both post-interaction trust measures also exhibited high reliability. The Trust Perception Scale-HRI (TPS-HRI) showed strong internal consistency overall ($\alpha$ = .89), with subscale reliabilities ranging from .74 to .88. The Trust in Industrial Human-Robot Collaboration scale (TI-HRC) demonstrated very high internal consistency ($\alpha$ = .95).
@@ 2.4.2 Objective Measures
Self-report measures were complemented by objective interaction metrics derived from system logs and manually coded dialogue transcripts. These measures captured the behavioural and process-level dynamics of collaboration, including interaction timing, task performance, speech recognition quality, and the quality of spoken-language exchange are aspects of the interaction that self-report alone would not fully characterise.
Coded metrics were derived from interaction logs and manually coded dialogue transcripts. For the purposes of dialogue coding, a turn was defined as a participant utterance and the robot’s subsequent response, treated as a single exchange unit. Table 3 outlines objective task metrics including the average number of dialogue turns, session duration, response time (robot dialogue latency), and the number of engaged versus frustrated responses the robot detected — all obtained from log data. Coded metrics from manual transcript analysis included the proportion of turns characterized by communication breakdowns — defined as turns in which the ASR pipeline returned fragmented or linguistically incoherent output that prevented the robot from extracting sufficient content to respond meaningfully. This definition excludes silent turns, in which no speech was detected within a given timeframe and the robot issued a condition-appropriate check-in; silence was treated as a distinct interaction event and coded separately. Other coded metrics included human and robot reasoning, robot helpful and unhelpful guidance, robot and human affective engagement, and other task-relevant human and robot contributions. See Appendix C for a full coding scheme.
@@ 2.5 Participants
A total of 29 participants were recruited from the Laurentian University community via word of mouth and the SONA recruitment system. Eligibility criteria required being 18 years or older, fluent in spoken and written English, and having normal or corrected-to-normal hearing and vision. Participants received a \$15 gift card as compensation for their time; some students additionally received partial credit for participating depending on their program of study. All procedures were approved by the Laurentian University Research Ethics Board (ROMEO File: #6021966).
Although English fluency was an eligibility requirement, in-person sessions revealed meaningful variability in participants’ functional spoken-language proficiency that was not captured in the online screening stage. In several sessions, the researcher observed difficulty in communication with participants prior to the robot interaction; suggesting that self-reported or screener-assessed fluency did not always reflect the level of spoken-language proficiency required for real-time autonomous speech recognition. Given the spoken-language demands of the task, the researcher documented these observations in session notes as a methodological record, in anticipation of potential downstream effects on interaction quality.
Subsequent review of interaction transcripts and system logs confirmed that a subset of sessions exhibited severe and sustained communication failure, characterized by fragmented or unintelligible ASR output and stalled dialogue. In these sessions, the robot was unable to extract sufficient linguistic content to establish a conversation, respond meaningfully to participant input, or support task progression. These interaction failures are described in detail in the Analytic Strategy and Results sections.
@@ 2.5.1 Randomization Check
Participants were initially assigned to the responsive or control condition by Qualtrics using an automated randomization procedure at the time of sign-up. However, because several scheduled participants did not attend their in-person session, replacement participants were assigned to the next available session slot rather than re-randomized. As a result, the final condition assignment should be considered semi-random rather than strictly randomized.